# MMDocIR: Benchmarking Multimodal Retrieval for Long Documents

### v2

@@ 6 Related Work
DocCVQA Tito et al. (2021) proposes extracting information from a document image collection. However, it is limited by its small question set (20 questions). While PDF-MVQA Ding et al. (2024) is tailored for multimodal retrieval in biomedical articles, it is annotated by GPT-3.5-turbo rather than experts. SciMMIR Wu et al. (2024) also investigates multimodal retrieval but only provides image-caption pairs, lacking user queries paired with the corresponding document pages. Ma et al. (2024a) introduces two relevant datasets, namely Wiki-SS and DocMatix-IR. Wiki-SS is derived from natural questions Kwiatkowski et al. (2019) , wherein evidence passages are screenshots of Wikipedia pages. However, natural questions are primarily designed for text retrieval, and the provided screenshots may not consistently capture the ground-truth evidence, as only the front page is considered. DocMatix-IR is constructed from the large-scale DocMatix Laurençon et al. (2024) dataset using filtering and hard negative mining. However, the questions are generated by Phi-3-small Abdin et al. (2024) rather than human experts, and are not de-contextualized for retrieval task. ViDoRe Faysse et al. (2024) is the most relevant benchmark to MMDocIR. It integrates multiple DocVQA datasets and provides new industrial documents. Upon a thorough examination of the 2,400 questions, we find that over 80% questions exhibit notable limitations in terms of their complexity, contextual clarity, and the absence of complete document corpora. Refer to Appendix F for detailed quantitative and qualitative analysis of ViDoRe.
@@ 7 Conclusion
In conclusion, multimodal document retrieval presents a complex challenge that requires both understanding and integrating diverse data modalities beyond plain text. To more effectively evaluate these capabilities, we introduce the MMDocIR benchmark, which features the innovative dual-task retrieval capabilities targeting page-level and layout-level document granularity. The MMDocIR includes a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 73,843 questions, serving as a valuable resource for both training and evaluation of multimodal document retrieval. Our comprehensive empirical studies show that visual-driven retrievers significantly outperform text-driven ones, underscoring the importance of visual information in improving retrieval performance. Future work can expand upon these findings by optimizing retrieval algorithms to enhance both accuracy and efficiency of multimodal document retrieval systems.
@@ Appendix A Supplementary Experimental Results
In this section, we provide the full results of page-level and layout-level retrievals, which supplement our main results discussion in Section 5.3 and Section 5.4 respectively.
Specifically, Table A extends the main page-level results shown in Table 4.2 with the results of text retrievers using OCR-text. Table A extends the main layout-level results shown in Table 5 with the results of (i) text retrievers using OCR-text and (ii) visual retrievers using hybrid inputs.
{tabu}
lll|cccccccccc|cc
@@
Recall@$k=1$
Textual Retrieval
OCR-text
DPR 3.4 7.2 1.2 11.3 3.0 9.8 8.2 24.7 30.9 26.3 12.6 12.4
@@
[…]
Visual Retrieval
Pure-Image
DSEwiki-ss 20.6 15.1 31.0 31.1 20.1 29.2 22.0 39.3 37.5 35.8 28.2 29.2
@@
Recall@$k=5$
Textual Retrieval
OCR-text
DPR 7.3 12.5 5.6 24.0 8.9 16.9 13.6 47.2 50.2 51.1 23.7 23.5
@@
[…]
Visual Retrieval
Pure-Image
DSEwiki-ss 42.4 32.9 56.3 58.5 39.8 50.6 41.6 68.6 60.9 50.0 50.2 52.1
@@
Recall@$k=10$
Textual Retrieval
OCR-text
DPR 10.5 21.1 8.8 32.0 14.9 19.6 17.0 59.7 56.4 58.9 29.9 29.3
@@
[…]
Visual Retrieval
Pure-Image
DSEwiki-ss 55.9 41.3 61.5 68.1 47.8 60.7 54.2 72.9 68.3 54.4 58.5 61.1
@@ B.1 Related DocVQA Benchmarks
Early DocVQA benchmarks primarily address single-page visual question answering (VQA), exemplified by datasets such as DocVQA Mathew et al. (2021), InfoVQA Mathew et al. (2022), and TAT-DQA Zhu et al. (2022). To overcome the limitations of single-page inputs, subsequent datasets like DUDE Landeghem et al. (2023), MP-DocVQA Tito et al. (2023), and SlideVQA Tanaka et al. (2023) have extended the context length to an average of 5.7, 8.3, and 20 pages, respectively. More recent benchmarks, including MMLongBench-Doc Ma et al. (2024b) and DocBench Zou et al. (2024), treat DocVQA as a long-context task, accommodating entire documents that average between 50–70 pages in length. As document lengths increase, retrieval becomes essential. Relevant pages must first be identified, followed by answer generation based on the retrieved multimodal evidence.
@@ B.2 Document Corpora Collection Criteria
To facilitate the development of MMDocIR, we leverage resources from existing DocVQA benchmarks described in Appendix B.1. Despite not being curated for IR, they offer valuable document corpora and questions that can be adapted for IR tasks. We select relevant DocVQA datasets based on the following criteria:
@@ B.4 Training Document Collection
  • DUDE Landeghem et al. (2023) provides 5,019 documents from aggregator websites7771: archive.org, 2: http://commons.wikimedia.org/, 3: http://documentcloud.org/. It covers a broad range of domains, including medical, legal, technical, and financial, among others, to evaluate models’ ability to handle diverse topics and the specific knowledge each requires. We filter out short documents and obtain 779 relatively long documents (averaging 15.6 pages) with 3,173 QA pairs.
@@ B.6 Hard Negative Sampling
In addition to annotating ground truth (positive) page labels, we enhance our training data with negative labels. In the context of retrieval, hard negatives are particularly informative non-relevant documents that closely resemble true positives according to the model’s current scoring function. Unlike randomly selected negatives, hard negatives are challenging for the model to distinguish from relevant documents, thus providing stronger supervision.
@@
As described in Appendix C.3, training is conducted using a contrastive loss, where the model aims to separate relevant documents from irrelevant ones. Specifically, we obtain hard negatives using the ColPali retriever Faysse et al. (2024), which scores all document pages for a given query. The irrelevant pages with the highest top-$k$ scores (i.e., those most likely to be confused with positives) are selected as hard negatives for training.
@@
MP-DocVQA Train-set curation Section 4 Appendix B.4 https://rrc.cvc.uab.es/?ch=17&com=tasks
@@
Text Retriever Layout location BGE and E5 (see Table D.1)
Visual Retriever Hard negatives Appendix B.6 Colpali (see Table D.1)
@@ C.2 Query-Doc Similarity
[…]
@@
[…]
@@ C.3 Contrastive Loss
($\mathcal{L}^{dpr}_{(q,d^{+},d^{-})}=-\log\frac{\exp(\text{Sim}^{dpr}_{(q,d^{+})}/\tau)}{\sum_{d_{i}\in d^{+}\cup d^{-}}\exp(\text{Sim}^{dpr}_{(q,d_{i})}/\tau)}$ (5)
@@ C.4 Training Implementation Details
In summary, we train two visual retrievers based on Phi3-Vision Abdin et al. (2024). DPR-Phi3 and Col-Phi3 represent document page or query using a single dense embedding (following DPR Karpukhin et al. (2020)) and a list of token-level embeddings (following ColBERT Khattab and Zaharia (2020)), respectively. To train the model, we employ memory-efficient techniques such as LoRA Hu et al. (2022), FlashAttention Dao (2024), and DeepSpeed Rasley et al. (2020).
The model is trained with a batch size of 64 for one epoch on MMDocIR training set. The model weights are shared between the language models for document screenshot and query encoding. In both tasks, each training query is paired with one positive document and one hard negative document. The document screenshots are resized to $344\times 1,$ 344 × 1 , 344 pixels and cropped into $4\times 4$ sub-images. The training process is performed on four A100 80GB GPUs.
@@ D.1 Text-Centric Document Retrieval
For text retrieval, the first step is to convert multimodal document into text using techniques, e.g., Document Parsing Chao and Fan (2004); Wang et al. (2024), Optical Character Recognition (OCR) Chaudhuri et al. (2017); Borovikov (2014); Mori et al. (1999), Layout Detection Sassioui et al. (2023); Xu et al. (2020, 2021), Chunking Chen et al. (2024a); Raina and Gales (2024); Dong et al. (2024), and Image Captioning You et al. (2016); Aneja et al. (2018). These steps are time-consuming and can introduce errors that impact the overall retrieval performance. Current text retrieval are primarily categorized as sparse or dense retrieval. For two widely-used sparse retrievers: TF-IDF Salton et al. (1983) calculates the relevance via word frequency with the inverse document frequency, and BM25 Robertson et al. (1994) introduces nonlinear word frequency saturation and length normalization. Dense retrievers encode content into vector representations. DPR Karpukhin et al. (2020) is the pioneering work of dense vector representations for QA tasks. Similarly, ColBERT Khattab and Zaharia (2020) introduces an efficient question-document interaction model with late fine-grained term matching. Contriever Izacard et al. (2022) leverages contrastive learning to improve content dense encoding. E5 Wang et al. (2022) and BGE Xiao et al. (2023) propose novel training and data preparation techniques to enhance retrieval performance. Moreover, GTE Li et al. (2023) integrates graph-based techniques to enhance dense embedding. However, most text retrieval systems overlook valuable visual information present in documents.
{tabu}
l|l|c|c|l Model Dimension Base Model HuggingFace Checkpoint
Text
DPR 768 BERT-base facebook/dpr-ctx_encoder-multiset-base facebook/dpr-question_encoder-multiset-base
[…]
@@
Visual
DSEwiki-ss 3,072 Phi-3-Vision Tevatron/dse-phi3-v1.0
[…]
@@
[…]
ColPali $N_{\mathrm{tok}}\times$1,024 PaliGemma vidore/colpali
@@ D.3 Implementation Details
In our experiments (refer to Section 5.2), we implement 9 off-the-shelf retrievers including 6 text retrievers and 3 visual retrievers. The text retrieval models deployed are namely DPR, ColBERT, Contriever, E5, BGE and GTE. These models use the WordPiece tokenizer from BERT and also inherit the maximum input length of 512 tokens from BERT Devlin et al. (2019). Additionally, we make use of the sentence-transformer library [^12] https://www.sbert.net/ when deploying E5, BGE and GTE. The visual retrieval models deployed are namely DSEwiki-ss, DSEdocmatix, and ColPali. We use pre-trained checkpoints available on HuggingFace  [^13] https://huggingface.co/; the specific checkpoint information can be found in Table D.1 alongside other configuration details.
@@ F.1 Query and Annotation Analysis
The comprehensive analysis of our queries, as presented in Table 14, reveals a significant challenge in adapting questions from Visual Question Answering (VQA) datasets (such as DocVQA, InfoVQA, TAT-DQA, and arXiVQA) for Information Retrieval (IR) purposes. Only 8% of these queries prove suitable for effective IR usage. In comparison, queries derived from industrial documents perform slightly better, with 15.5% deemed suitable. A common issue identified is that these queries are either excessively simplistic or highly specific to a particular context. Our findings suggest that the primary difficulty stems from the inherent differences between DocIR and DocVQA. VQA queries are typically crafted to address content on a specific page or within a particular image, inherently limiting their scope and specificity. This specificity and simplism are functional within the confines of the intended VQA context but pose substantial limitations when such queries are repurposed for IR tasks. Due to this gap, we exclude ViDoRe benchmark from our experiments.

### v3

@@ 6 Related Work
DocCVQA Tito et al. (2021) proposes extracting information from a document image collection. However, it is limited by its small question set (20 questions). While PDF-MVQA Ding et al. (2024) is tailored for multimodal retrieval in biomedical articles, it is annotated by GPT-3.5-turbo rather than experts. SciMMIR Wu et al. (2024) also investigates multimodal retrieval but only provides image-caption pairs, lacking user queries paired with the corresponding document pages. Ma et al. (2024a) introduces two relevant datasets, namely Wiki-SS and DocMatix-IR. Wiki-SS is derived from natural questions Kwiatkowski et al. (2019) , wherein evidence passages are screenshots of Wikipedia pages. However, natural questions are primarily designed for text retrieval, and the provided screenshots may not consistently capture the ground-truth evidence, as only the front page is considered. DocMatix-IR is constructed from the large-scale DocMatix Laurençon et al. (2024) dataset using filtering and hard negative mining. However, the questions are generated by Phi-3-small Abdin et al. (2024) rather than human experts, and are not de-contextualized for retrieval task. MMDocRAG Dong et al. (2025a) is constructed upon MMDocIR to support multimodal generation. ViDoRe Faysse et al. (2024) is the most relevant benchmark to MMDocIR. It integrates multiple DocVQA datasets and provides new industrial documents. Upon a thorough examination of the 2,400 questions, we find that over 80% questions exhibit notable limitations in terms of their complexity, contextual clarity, and the absence of complete document corpora. Refer to Appendix F for detailed quantitative and qualitative analysis of ViDoRe.
@@ 7 Conclusion
In conclusion, multimodal document retrieval presents a complex challenge that requires both understanding and integrating diverse data modalities beyond plain text. To more effectively evaluate these capabilities, we introduce the MMDocIR benchmark, which features the innovative dual‐task retrieval capabilities targeting page-level and layout-level document granularity. The MMDocIR includes a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 73,843 questions, serving as a valuable resource for both training and evaluation of multimodal document retrieval. Our comprehensive empirical studies show that visual-driven retrievers significantly outperform text-driven ones, underscoring the importance of visual information in improving retrieval performance. Future work can expand upon these findings by optimizing retrieval algorithms to enhance both accuracy and efficiency of multimodal document retrieval systems, as well as the multilingual capability Liang et al. (2020).
@@ Appendix A Supplementary Experimental Results
In this section, we provide the full results of page-level and layout-level retrievals, which supplement our main results discussion in Section 5.3 and Section 7 respectively.
Specifically, Table 11 extends the main page-level results shown in Table 5 with the results of text retrievers using OCR-text. Table 12 extends the main layout-level results shown in Table 6 with the results of (i) text retrievers using OCR-text and (ii) visual retrievers using hybrid inputs.
@@
Recall@$k=1$ Textual Retrieval OCR-text DPR 3.4 7.2 1.2 11.3 3.0 9.8 8.2 24.7 30.9 26.3 12.6 12.4
@@
Visual Retrieval Pure-Image DSEwiki-$s$ 20.6 15.1 31.0 31.1 20.1 29.2 22.0 39.3 37.5 35.8 28.2 29.2
@@
Recall@$k=5$ Textual Retrieval OCR-text DPR 7.3 12.5 5.6 24.0 8.9 16.9 13.6 47.2 50.2 51.1 23.7 23.5
@@
Visual Retrieval Pure-Image DSEwiki-$s$ 42.4 32.9 56.3 58.5 39.8 50.6 41.6 68.6 60.9 50.0 50.2 52.1
@@
Recall@$k=10$ Textual Retrieval OCR-text DPR 10.5 21.1 8.8 32.0 14.9 19.6 17.0 59.7 56.4 58.9 29.9 29.3
@@
Visual Retrieval Pure-Image DSEwiki-$s$ 55.9 41.3 61.5 68.1 47.8 60.7 54.2 72.9 68.3 54.4 58.5 61.1
@@ B.1 Related DocVQA Benchmarks
Early DocVQA benchmarks primarily address single-page visual question answering (VQA), exemplified by datasets such as DocVQA Mathew et al. (2021), InfoVQA Mathew et al. (2022), and TAT-DQA Zhu et al. (2022). To overcome the limitations of single-page inputs, subsequent datasets like DUDE Landeghem et al. (2023), MP-DocVQA Tito et al. (2023), and SlideVQA Tanaka et al. (2023) have extended the context length to an average of 5.7, 8.3, and 20 pages, respectively. More recent benchmarks, including MMLongBench-Doc Ma et al. (2024b) and DocBench Zou et al. (2024), treat DocVQA as a long-context task, accommodating entire documents that average between 50 to 70 pages in length. As document lengths increase, retrieval becomes essential. Relevant pages must first be identified, followed by answer generation based on the retrieved multimodal evidence.
@@ B.2 Document Corpora Collection Criteria
Early document related benchmarks Dong et al. (2021) are mostly textual only, which are not considered in MMDocIR. To facilitate the development of MMDocIR, we leverage visually-rich documents from recent DocVQA benchmarks described in Appendix B.1. Despite not being curated for IR, they offer valuable document corpora and questions that can be adapted for IR tasks. We select relevant DocVQA datasets based on the following criteria:
@@ B.4 Training Document Collection
  • DUDE Landeghem et al. (2023) provides 5,019 documents from aggregator websites1212121: archive.org, 2: http://commons.wikimedia.org/, 3: http://documentcloud.org/. It covers a broad range of domains, including medical, legal, technical, and financial, among others, to evaluate models’ ability to handle diverse topics and the specific knowledge each requires. We filter out short documents and obtain 779 relatively long documents (averaging 15.6 pages) with 3,173 QA pairs.
@@ B.6 Hard Negative Sampling
In addition to annotating ground truth (positive) page labels, we enhance our training data with negative labels Li et al. (2025). In the context of retrieval, hard negatives are particularly informative non-relevant documents that closely resemble true positives according to the model’s current scoring function. Unlike randomly selected negatives, hard negatives are challenging for the model to distinguish from relevant documents, thus providing stronger supervision.
@@
As described in Appendix C.3, training is conducted using a contrastive loss, where the model aims to separate relevant documents from irrelevant ones. Specifically, we obtain hard negatives using the ColPali retriever Faysse et al. (2024), which scores all document pages for a given query. The irrelevant pages with the highest top-$k$ scores (i.e., those most likely to be confused with positives) are selected as hard negatives for training. In the future, we consider to incorporate content planning Bao et al. (2022b) and synthetic methods Bao et al. (2023, 2022a) for hard negative generation.
@@
MP-DocVQA Train-set curation
Section 4
Appendix B.4
https://rrc.cvc.uab.es/?ch=17&com=tasks
@@
Text Retriever Layout location BGE and E5 (see Table 14)
Visual Retriever Hard negatives Appendix B.6 Colpali (see Table 14)
@@ C.2 Query-Doc Similarity
where $Sim(q,d)_{dpr}$ is computed as the cosine similarity between their embeddings. and $\langle\cdot|\cdot\rangle$ is the dot product.
@@
where $Sim(q,d)_{col}$ is the sum over all query vectors $\mathrm{E_{q}^{col}}^{(i)}$, of its maximum dot product $\langle\cdot|\cdot\rangle$ with each of the $N_{d}$ document embedding vectors $\mathrm{E_{d}^{col}}^{(j)}$.
@@ C.3 Contrastive Loss
$\mathcal{L}^{dpr}_{(q,d^{+},d^{-})}=-\log\frac{\exp(\text{Sim}^{dpr}_{(q,d^{+})}/\tau)}{\sum_{d_{i}\in d^{+}\cup d^{-}}\exp(\text{Sim}^{dpr}_{(q,d_{i})}/\tau)}$ (5)
@@ C.4 Training Implementation Details
In summary, we train two visual retrievers based on Phi3-Vision Abdin et al. (2024). DPR-Phi3 and Col-Phi3 represent document page or query using a single dense embedding (following DPR Karpukhin et al. (2020)) and a list of token-level embeddings (following ColBERT Khattab and Zaharia (2020)), respectively. To train the model, we employ memory-efficient techniques such as PERF Zhang et al. (2024b, a), LoRA Hu et al. (2022), FlashAttention Dao (2024), and DeepSpeed Rasley et al. (2020).
The model is trained with a batch size of 64 for one epoch on MMDocIR training set. The model weights are shared between the language models for document screenshot and query encoding. In both tasks, each training query is paired with one positive document and one hard negative document. The document screenshots are resized to $1,344\times 1,344$ pixels and cropped into $4\times 4$ sub-images.
@@ D.1 Text-Centric Document Retrieval
For text retrieval, the first step is to convert multimodal document into text using techniques, e.g., Document Parsing Chao and Fan (2004); Wang et al. (2024), Optical Character Recognition (OCR) Chaudhuri et al. (2017); Borovikov (2014); Mori et al. (1999), Layout Detection Sassioui et al. (2023); Xu et al. (2020, 2021), Information extraction Dong et al. (2022, 2023a), Chunking Chen et al. (2024a); Raina and Gales (2024), and Image Captioning You et al. (2016); Aneja et al. (2018). These steps are time-consuming and can introduce errors that impact the overall retrieval performance Wu et al. (2025); Nie et al. (2023); Li et al. (2023a). Current text retrieval are primarily categorized as sparse or dense retrieval on chunks Dong et al. (2023b). For two widely-used sparse retrievers: TF-IDF Salton et al. (1983) calculates the relevance via word frequency with the inverse document frequency, and BM25 Robertson et al. (1994) introduces nonlinear word frequency saturation and length normalization. Dense retrievers encode content into vector representations. DPR Karpukhin et al. (2020) is the pioneering work of dense vector representations for QA tasks. Similarly, ColBERT Khattab and Zaharia (2020) introduces an efficient question-document interaction model with late fine-grained term matching. Contriever Izacard et al. (2022) leverages contrastive learning to improve content dense encoding. E5 Wang et al. (2022) and BGE Xiao et al. (2023) propose novel training and data preparation techniques to enhance retrieval performance. Moreover, GTE Li et al. (2023b) integrates graph-based techniques to enhance dense embedding. However, most text retrieval systems overlook valuable visual information present in documents.
Model Dimension Base Model HuggingFace Checkpoint
Text DPR 768 BERT-base
facebook/dpr-ctx_encoder-multiset-base
facebook/dpr-question_encoder-multiset-base
@@
Visual DSEwiki-$s$ 3,072 Phi-3-Vision Tevatron/dse-phi3-v1.0
@@
ColPali $N_{\mathrm{tok}}\times$128 PaliGemma vidore/colpali
@@ D.3 Implementation Details
In our experiments (refer to Section 5.2), we implement 9 off-the-shelf retrievers including 6 text retrievers and 3 visual retrievers. The text retrieval models deployed are namely DPR, ColBERT, Contriever, E5, BGE and GTE. These models use the WordPiece tokenizer from BERT and also inherit the maximum input length of 512 tokens from BERT Devlin et al. (2019). Additionally, we make use of the sentence-transformer library [^17] https://www.sbert.net/ when deploying E5, BGE and GTE. The visual retrieval models deployed are namely DSEwiki-ss, DSEdocmatix, and ColPali. We use pre-trained checkpoints available on HuggingFace  [^18] https://huggingface.co/; the specific checkpoint information can be found in Table 14 alongside other configuration details.
@@ F.1 Query and Annotation Analysis
The comprehensive analysis of our queries, as presented in Table 15, reveals a significant challenge in adapting questions from Visual Question Answering (VQA) datasets (such as DocVQA, InfoVQA, TAT-DQA, and arXiVQA) for Information Retrieval (IR) purposes. Only 8% of these queries prove suitable for effective IR usage. In comparison, queries derived from industrial documents perform slightly better, with 15.5% deemed suitable. A common issue identified is that these queries are either excessively simplistic or highly specific to a particular context. Our findings suggest that the primary difficulty stems from the inherent differences between DocIR and DocVQA. VQA queries are typically crafted to address content on a specific page or within a particular image, inherently limiting their scope and specificity. This specificity and simplism are functional within the confines of the intended VQA context but pose substantial limitations when such queries are repurposed for IR tasks.